OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches zero of quantity x cubed plus one divided by x to the fifth power.

OpenStudy (anonymous):

please help!!

OpenStudy (amoodarya):

\[\lim_{x \rightarrow 0}\frac{ x^3+1 }{ x^5 }\] did you mean it ?

OpenStudy (anonymous):

yes

OpenStudy (amoodarya):

\[\lim_{x \rightarrow 0^+}\frac{ x^3+1 }{ x^5 }=\frac{ 1 }{ 0^+ }=+ \infty\\lim_{x \rightarrow 0^-}\frac{ x^3+1 }{ x^5 }=\frac{ 0^-+1 }{ 0^- }=+ \infty\]

OpenStudy (amoodarya):

sorry second was -inf

OpenStudy (anonymous):

okay so let me show you the answer choices

OpenStudy (amoodarya):

\[\lim_{x \rightarrow 0^-}\frac{ x^3+1 }{ x^5 }=\frac{ 0^-1+1 }{ 0^- }=\frac{ 1 }{ 0^- }=- \infty\]

OpenStudy (anonymous):

i think i know which one it might be

OpenStudy (anonymous):

0 -9 Does not exist 9

OpenStudy (amoodarya):

Does not exist

OpenStudy (anonymous):

okay i figured that, thanks so much!!

OpenStudy (anonymous):

do you have time for one more

OpenStudy (amoodarya):

write it

OpenStudy (anonymous):

Find the derivative of f(x) = negative 9 divided by x at x = -8

OpenStudy (amoodarya):

\[f(x)=\frac{ -9 }{ x }=?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

answer choices: 8 divided by 9 9 divided by 8 64 divided by 9 9 divided by 64

OpenStudy (amoodarya):

\[f(x)=\frac{ -9 }{ x }\\f'(x)=\frac{ 0(x)-(1)(-9) }{x^2 }\\f'=\frac{ 9 }{ x^2 }\\now\\put\\x=-8\]

OpenStudy (anonymous):

-64?